# Robust parameter estimation within minutes on gravitational wave signals from binary neutron star inspirals

### v1

@@ V.1 Science cases
The robustness and speed-up offered by Jim are crucial to meet the computational requirements of challenging science studies. For instance, while sky maps can be produced in low latency with bayestar Singer and Price (2016), performing PE on the intrinsic parameters improves the knowledge of the source properties and would be essential to improve estimates of electromagnetic (EM) emission, such as the brightness of a potential kilonova; see, e.g., Ref. Stachie et al. (2021). In general, good guidance for EM follow-up searches is critical for improving the scientific return of follow-up campaigns while on the other hand lowering the requirement for expensive telescope time. As such, the speed of Jim in performing PE offers a way to enhance low-latency follow-up strategies.
Furthermore, the next generation of GW detectors, such as Einstein Telescope Punturo et al. (2010), will have an overall greater sensitivity and broader frequency bandwidth compared to existing detectors Maggiore et al. (2020); Evans et al. (2023); Branchesi et al. (2023). Future detectors will observe more BNS inspirals, which will also be in the sensitive band for a longer time, calling for efficient and effective PE pipelines Samajdar et al. (2021); Pizzati et al. (2022).
@@ V.2 Comparison to related works
One likelihood-based method to speed up PE is relative binning and was discussed in Sec.II.3. Previous works have used relative binning for multi-messenger and low-latency PE studies Finstad and Brown (2020); Raaijmakers et al. (2021). Relative binning was also recently integrated into Bilby Krishna et al. (2023).
Another way to speed up the likelihood evaluation is through reduced order modelling (ROM) and ROQ methods. Using ROQ, previous works have been able to perform PE in the order of minutes Canizares et al. (2015); Morisaki and Raymond (2020); Morisaki et al. (2023). However, ROQ requires precomputing these reduced bases, which can be computationally expensive, as we show with our estimate in Appendix C. While these bases have to be constructed only once, Jim, on the other hand, does not require any precomputed quantities as input for the PE.
Besides relative binning and ROQ, other approximations of the likelihood have been studied as well, such as approximating the likelihood with Gaussian processes Pankow et al. (2015); Lange et al. (2018); Wysocki et al. (2019), a mesh-free interpolation method Pathak et al. (2023, 2024), or through marginalization of the likelihood, e.g.  Roulet et al. (2024).
Apart from approximating the likelihood, ML can speed up sampling as well. For example, nessai accelerates nested sampling through normalizing flows Williams et al. (2021). Recently, this method achieved wall times similar to ours by combining nessai with the ROQ approximation Williams et al. (2023). While similar in spirit to our work, we note that Jim makes use of MCMC rather than nested sampling for exploring the likelihood landscape, and uses relative binning rather than ROQ to speed up likelihood evaluation, thereby removing the need to precompute the ROQ bases.
Finally, likelihood-free methods to accelerate PE mainly consist of ML models pretrained on simulated data to approximate the likelihood or posterior distributions Gabbard et al. (2022); Kolmus et al. (2022, 2024); Chua and Vallisneri (2020); Green et al. (2020); Green and Gair (2021); Dax et al. (2021, 2023); Bhardwaj et al. (2023). Contrary to these pipelines, Jim conducts the PE without the need for pretraining.
@@ V.3 Environmental impact
We compare the carbon footprint of Jim relative to the other pipelines reported in Tab. 4. The TDP of an NVIDIA A100-$40$ GB GPU, used for the Jim runs, is $400$NVIDIA Corporation (2024). The pBilby runs used a single Intel Xeon Platinum 8174 Processor with a TDP of $240$Intel Corporation (2024a). The RB-Bilby and ROQ-Bilby used a single Intel Xeon Silver 4310 Processor CPU with a TDP of $120$Intel Corporation .
Based on our PE runs, we take the average wall time for each pipeline, and the TDP reported above to estimate the required energy to produce the results for all $200$ injection runs and the $4$ real event runs shown in this work. The results are reported in amounts of kWh in Tab. 5.777The obtained numbers have been rounded for clarity and ease of presentation. In order to make a fair comparison against Jim, we also report the time taken up by the pretraining phase of the ROQ method, i.e., the time taken to construct the ROQ bases. The details of this estimate can be found in Appendix C. For comparison, the average annual electricity consumption of a Netherlands household in 2021 was 28102810{2810}2810 kWh Statistics Netherlands (2023).
To translate these into tangible numbers, we will assume that $0.328$ kg $\rm{CO}_{2}$ is produced per kWh CO2 emissiefactoren (2024) and that it approximately takes $50$ trees a year to capture 10001000{1000}1000 kg of $\rm{CO}_{2}$ Anthesis Climate Neutral group (2024). Given a year, it takes around $0.55\rm{\ trees}$ to absorb the amount of $\rm{CO}_{2}$ generated by Jim. On the other hand, around $59.02$ and $1.49\rm{\ trees}$ are needed for absorbing the $\rm{CO}_{2}$ generated by pBilby and RB-Bilby, respectively. While only $0.52\rm{\ trees}$ are needed to counter the carbon footprint from the sampling of ROQ-Bilby, the precompute of the ROQ bases requires around $0.44\rm{\ trees}$. We can, therefore, conclude that Jim is more environmentally friendly than the other pipelines considered in performing the PE runs in this work. However, we would like to emphasize that the cost associated with building the ROQ bases is a one-time cost only. Therefore, we estimate that ROQ-Bilby will have a similar ecological footprint as Jim after around $3000$ PE runs.
kWh $\rm{CO}_{2}$ [kg] Trees
[…]
pBilby $3599$ $1180$ $59.02$
RB-Bilby $91$ $30$ $1.49$
@@ V.4 Future work
While our current approach, as outlined in Sec.II.3, allows us to use IMRPhenomD_NRTidalv2 with the existing relative binning implementation, it is desirable to extend the relative binning method to waveforms which include a tapering window, such as the NRTidal family of waveforms. For this, one requires a scheme to construct the bins without making prior assumptions regarding the waveforms used. An example of such an agnostic approach is given by Ref. Narola et al. (2023).
In addition, a current bottleneck in the development of Jim is the need for a jax-compatible implementation of existing GW approximants. At the time of writing, a ripple implementation of the precessing waveforms IMRPhenomPv2 and IMRPhenomXPHM are under development. Additionally, we will investigate the possibility of training surrogate ML models to approximate the waveforms, which removes the need to re-implement waveforms from scratch and reduces development time.
Furthermore, future work can investigate the integration with existing packages. While our current pipeline produces posterior samples of any desired target distribution, it does not output the evidence, which is crucial for model selection. In the future, we wish to estimate the Bayesian evidence by performing importance sampling on our produced samples. Future work can integrate Jim with denmarf Lo (2023) or harmonic McEwen et al. (2023); Spurio Mancini et al. (2022); Polanska et al. (2024) to obtain the Bayesian evidence. Secondly, one can attempt to integrate the recently introduced package astreos McGinn et al. (2024) with our pipeline. astreos can generate confidence intervals for the equations of state in less than one second by relying on normalizing flows. However, the NF maps the component masses and tidal deformabilities to a set of so-called auxiliary parameters from which the EOS constraints can be inferred. Obtaining estimates for these intrinsic parameters can be done reliably and efficiently with Jim, which can therefore empower astreos.
@@ VI Conclusions
In this work, we have demonstrated the robustness, accuracy, and speed of Jim in performing parameter estimation for gravitational waves of binary neutron star mergers with tidal effects. By combining relative binning, jax, gradient-based MCMC and normalizing flows, Jim can produce posteriors accurately in the order of minutes without requiring any precomputed input such as a pretrained normalizing flow. Specifically, we are able to analyze GW170817 and GW190425 with the TaylorF2 and IMRPhenomD_NRTidalv2 waveforms in $21$ to $33$ minutes of total wall time, depending on the sampling settings and the number of bins used in relative binning. Evaluated on a large suite of simulated events, the median total wall times are below $30$ minutes for both waveforms. As such, Jim can be an indispensable tool for future science cases, such as low-latency follow-up strategies and data analysis with next-generation gravitational wave detectors. Additionally, we have shown that Jim has significantly reduced the environmental impact of parameter estimation as compared to other pipelines. Therefore, Jim presents an ecologically sustainable approach to addressing these computationally intensive tasks.
@@ Appendix C Estimate of ROQ environmental impact
We provide details on how the estimate of the environmental impact of constructing ROQ bases, reported in Tab. 5, is computed based on Ref. Soichiro Morisaki (2024).
Building the ROQ bases for the IMRPhenomPv2_NRTidalv2 waveform takes around $\sim 1$ day when using $\sim 100$ cores on the NEMO cluster owned by UWM when making use of multi-banding Morisaki (2021) to speed up the building of the ROQ bases. We will assume around $\sim 5$ Intel Xeon Gold 6136 Processor CPUs were used, each having a TDP of $150$Intel Corporation (2024b). Therefore, we estimate the energy consumption of building the ROQ bases for this waveform to be around $27$ kWh.
@@
We would like to emphasize that this is only an estimate of the computational cost. In particular, further advancements in the techniques used for building ROQ bases can potentially reduce the computational cost, e.g. Ref. Morras et al. (2023).

### v2

@@ V.1 Science cases
The robustness and speed-up offered by Jim are crucial to meet the computational requirements of challenging science studies. For instance, while sky maps can be produced in low latency with bayestar Singer:2015ema, performing PE on the intrinsic parameters improves the knowledge of the source properties and would be essential to improve estimates of electromagnetic (EM) emission, such as the brightness of a potential kilonova; see, e.g., Ref. Stachie:2021noh. In general, good guidance for EM follow-up searches is critical for improving the scientific return of follow-up campaigns while on the other hand lowering the requirement for expensive telescope time. As such, the speed of Jim in performing PE offers a way to enhance low-latency follow-up strategies.
Furthermore, the next generation of GW detectors, such as Einstein Telescope Punturo:2010zza and Cosmic Explorer Evans:2021gyd, will have an overall greater sensitivity and broader frequency bandwidth compared to existing detectors Maggiore:2019uih; Evans:2023euw; Branchesi:2023mws. Future detectors will observe more BNS inspirals, which will also be in the sensitive band for a longer time, calling for efficient and effective PE pipelines Samajdar:2021egv; Pizzati:2021apa.
@@ V.2 Comparison to related works
One likelihood-based method to speed up PE is relative binning and was discussed in Sec. II.3. Previous works have used relative binning for multi-messenger and low-latency PE studies Finstad:2020sok; Raaijmakers:2021slr. Relative binning was also recently integrated into Bilby Krishna:2023bug.
Another way to speed up the likelihood evaluation is through reduced order modelling (ROM) and ROQ methods. Using ROQ, previous works have been able to perform PE in the order of minutes Canizares:2014fya; Morisaki:2020oqk; Morisaki:2023kuq. However, ROQ requires precomputing these reduced bases, which can be computationally expensive, as we show with our estimate in Appendix C. While these bases have to be constructed only once, Jim, on the other hand, does not require any precomputed quantities as input for the PE.
Besides relative binning and ROQ, other approximations of the likelihood have been studied as well, such as approximating the likelihood with Gaussian processes Pankow:2015cra; Lange:2018pyp; Wysocki:2019grj and a mesh-free interpolation method Pathak:2022iar; Pathak:2023ixb. Alternatively, one can marginalize over certain parameters in the likelihood to make parameter estimation more efficient and robust, such as in Ref. Roulet:2024hwz.
Apart from speeding up the evaluation of the likelihood, ML can be used to improve the efficiency of the sampler and speed up PE. For example, nessai accelerates nested sampling through normalizing flows Williams:2021qyt. Recently, this method achieved wall times similar to ours by combining nessai with the ROQ approximation Williams:2023ppp. While similar in spirit to our work, we note that Jim makes use of MCMC rather than nested sampling for exploring the likelihood landscape, and uses relative binning rather than ROQ to speed up likelihood evaluation, thereby removing the need to precompute the ROQ bases.
Finally, likelihood-free methods to accelerate PE mainly consist of ML models pretrained on simulated data to approximate the likelihood or posterior distributions Gabbard:2019rde; Kolmus:2021buf; Kolmus:2024scm; Chua:2019wwt; Green:2020hst; Green:2020dnx; Dax:2021tsq; Dax:2022pxd; Bhardwaj:2023xph. Contrary to these pipelines, Jim conducts the PE without the need for pretraining.
@@ V.3 Environmental impact
We compare the carbon footprint of Jim relative to the other pipelines reported in Tab. 6. The TDP of an NVIDIA A100-$40$ GB GPU, used for the Jim runs, is 400400nvidia_a100_datasheet. The pBilby runs used a single Intel Xeon Platinum 8174 Processor with a TDP of 240240ark_intel_cpu_datasheet. The RB-Bilby and ROQ-Bilby used a single Intel Xeon Silver 4310 Processor CPU with a TDP of 120120intel_cpu_datasheet_4310.
Based on our PE runs, we take the average wall time for each pipeline, and the TDP reported above to estimate the required energy to produce the results for all $200$ injection runs and the $4$ real event runs shown in this work. The results are reported in amounts of kWh in Tab. 6.777The obtained numbers have been rounded for clarity and ease of presentation. In order to make a fair comparison against Jim, we also report the time taken up by the pretraining phase of the ROQ method, i.e., the time taken to construct the ROQ bases. The details of this estimate can be found in Appendix C. For comparison, the average annual electricity consumption of a Netherlands household in 2021 was 2810{2810} kWh Statistics_Netherlands_2023.
To translate these into tangible numbers, we will assume that $0.328$ kg $\rm{CO}_{2}$ is produced per kWh NL_CO2_emissiefactoren and that it approximately takes $50$ trees a year to capture 1000{1000} kg of $\rm{CO}_{2}$ climateneutral_trees_reference. Given a year, it takes around $0.55\rm{\ trees}$ to absorb the amount of $\rm{CO}_{2}$ generated by Jim. On the other hand, around $67.68$ and $1.32\rm{\ trees}$ are needed for absorbing the $\rm{CO}_{2}$ generated by pBilby and RB-Bilby, respectively. While only $0.52\rm{\ trees}$ are needed to counter the carbon footprint from the sampling of ROQ-Bilby, the precompute of the ROQ bases requires around $0.44\rm{\ trees}$. We can, therefore, conclude that Jim is more environmentally friendly than the other pipelines considered in performing the PE runs in this work. However, we would like to emphasize that the cost associated with building the ROQ bases is a one-time cost only. Therefore, we estimate that ROQ-Bilby will have a similar ecological footprint as Jim after around $3000$ PE runs.
kWh $\rm{CO}_{2}$ $10^{3}$ kg] Trees
Jim $\phantom{00}34$ $\phantom{0}11$ $\phantom{000}0.55$
pBilby $4127$ $1354$ $67.68$
RB-Bilby $80$ $26$ $1.32$
@@ V.4 Future work
While our current approach, as outlined in Sec. II.3, allows us to use IMRPhenomD_NRTidalv2 with the existing relative binning implementation, it is desirable to extend the relative binning method to waveforms which include a tapering window, such as the NRTidal family of waveforms. For this, one requires a scheme to construct the bins without making prior assumptions regarding the waveforms used. An example of such an agnostic approach is given by Ref. Narola:2023men.
In addition, a current bottleneck in the development of Jim is the need for a jax-compatible implementation of existing GW approximants. At the time of writing, a ripple implementation of the precessing waveforms IMRPhenomPv2 and IMRPhenomXPHM are under development. Recently, a jax-compatible implementation of the IMRPhenomPv2_NRTidal approximant has been constructed Dax:2024mcn. Additionally, we will investigate the possibility of training surrogate ML models to approximate the waveforms, which removes the need to re-implement waveforms from scratch and reduces development time.
Furthermore, future work can investigate the integration with existing packages. While our current pipeline produces posterior samples of any desired target distribution, it does not output the evidence, which is crucial for model selection. In the future, we wish to estimate the Bayesian evidence by performing importance sampling on our produced samples. Future work can integrate Jim with denmarf lo2023denmarf or harmonic mcewen2023machine; SpurioMancini:2022vcy; polanska2024learned to obtain the Bayesian evidence. Secondly, one can attempt to integrate the recently introduced package astreos McGinn:2024nkd with our pipeline. astreos can generate confidence intervals for the equations of state in less than one second by relying on normalizing flows. However, the NF maps the component masses and tidal deformabilities to a set of so-called auxiliary parameters from which the EOS constraints can be inferred. Obtaining estimates for these intrinsic parameters can be done reliably and efficiently with Jim, which can therefore empower astreos.
@@ VI Conclusions
In this work, we have demonstrated the robustness, accuracy, and speed of Jim in performing parameter estimation for gravitational waves of binary neutron star mergers with tidal effects. By combining relative binning, jax, gradient-based MCMC and normalizing flows, Jim can produce posteriors accurately in the order of minutes without requiring any precomputed input such as a pretrained normalizing flow. Specifically, we are able to analyze GW170817 and GW190425 with the TaylorF2 and IMRPhenomD_NRTidalv2 waveforms in $20$ to $31$ minutes of total wall time, depending on the sampling settings and the number of bins used in relative binning. Evaluated on a large suite of simulated events, the median total wall times are below $30$ minutes for both waveforms. As such, Jim can be an indispensable tool for future science cases, such as low-latency follow-up strategies and data analysis with next-generation gravitational wave detectors. Additionally, we have shown that Jim has significantly reduced the environmental impact of parameter estimation as compared to other pipelines. Therefore, Jim presents an ecologically sustainable approach to addressing these computationally intensive tasks.
@@ Appendix C Estimate of ROQ environmental impact
We provide details on how the estimate of the environmental impact of constructing ROQ bases, reported in Tab. 6, is computed based on Ref. Soichiro_mail.
Building the ROQ bases for the IMRPhenomPv2_NRTidalv2 waveform takes around $\sim 1$ day when using $\sim 100$ cores on the NEMO cluster owned by UWM when making use of multi-banding Morisaki:2021ngj to speed up the building of the ROQ bases. We will assume around $\sim 5$ Intel Xeon Gold 6136 Processor CPUs were used, each having a TDP of 150150intel_cpu_datasheet_gold_8380. Therefore, we estimate the energy consumption of building the ROQ bases for this waveform to be around $27$ kWh.
@@
We would like to emphasize that this is only an estimate of the computational cost. In particular, further advancements in the techniques used for building ROQ bases can potentially reduce the computational cost, e.g. Ref. Morras:2023pug.
@@ Acknowledgments
ESS
effective sample size
RNG
random number generator